# Simulating Chemistry on Bosonic Quantum Devices

### v1

@@ 1 Introduction
In this paper, we focus on a promising experimental realization of a qumode-based platform that uses circuit quantum electrodynamics (cQED). 15 In this case, superconducting resonators serve as qumodes, with the interactions between resonators controlled via an auxiliary superconducting circuit. 16, 17, 18 Other hardware approaches to bosonic quantum computing include arrays of trapped ions, 19 optical modes, 20, or mechanical resonators. 21
The remainder of this perspective is organized as follows. An introduction to bosonic quantum computing is provided in Section 2. Recent progress in simulating molecular vibronic spectra22, 23 and chemical quantum dynamics 24 on bosonic quantum devices is reviewed in Section 3. The potential for bosonic quantum devices to solve a wide range of timely problems in computational chemistry, including nonadiabatic chemical dynamics, the efficient solution of molecular graph theory problems and electronic structure calculations is highlighted in Section 4. Concluding remarks are provided in Section 5.
@@ 2.1 A primer on bosonic systems
where {|$\{\ket{q}|\>q\in\mathbb{R}\}$ and {|$\{\ket{p}|\>p\in\mathbb{R}\}$ are two possible continuous variable (CV) bases for a single qumode Hilbert space. The transformation between the position and momentum bases is given by the Fourier transform 25
@@
The existence of position and momentum CV bases also makes it possible to represent quantum states in terms of wave functions $\langle q|\psi\rangle=\psi(q)$ and $\langle p|\psi\rangle=\tilde{\psi}(p)$, in the position or momentum representations, respectively. Of particular importance for the bosonic devices under consideration in this paper are Gaussian states, a class of CV states that can be represented in terms of Gaussian functions. 25 One such Gaussian state is the ground state (also known as the vacuum state) of a single qumode, $\langle q|0\rangle=\psi_{0}(q)$. Thus, Gaussian transformations such as phase space displacement and squeezing operators can be defined as operators that map Gaussian states to Gaussian states. 25 Some of the Gaussian unitary operators are listed in Table 1. However, for universal quantum computations, the Gaussian operations have to be combined with non-Gaussian unitaries. 26, 27 The implementation of non-Gaussian operations can be efficiently done on the bosonic circuit quantum electrodynamics (cQED) devices, as discussed below.
@@ 2.2 Hardware
Each resonator can be thought of as a quantum LC circuit, consisting of an inductor and a capacitor, with the resonant frequency $\omega=1/\sqrt{LC}$ depending on inductance $L$ and capacitance $C$. The resonator geometry design critically affects the performance of the quantum devices. Early resonator prototypes focused on the planar design because of the convenience of fabricating superconducting metal circuits on the same chip that contains the auxiliary circuits with standard lithographic tools. 28 A range of geometries have been introduced since, including 3D resonators that can produce coherence time up to milliseconds by taking advantage of their hollow structure. 15 The coupling to the auxiliary circuit can now be achieved, for example, by introducing a center pin in a cylindrical resonator, while placing the auxiliary circuit next to the resonator. 18
The auxiliary circuits that are used to couple the qumodes are made of a Josephson junction, a device of two layers of superconducting material separated by a weak link such as a thin insulating tunnel barrier. The Josephson junction becomes insensitive to charge noise in the regime of large Josephson energy to the charging energy ratio, 29 which is commonly referred to as a transmon. 30, 31 The transmon can be understood as an anharmonic resonator and is coupled to the qumode resonator by extending its superconducting layers to overlap with the resonator’s electromagnetic field. The interaction between the auxiliary circuits and qumodes can be controlled by detuning the qumode and transmon resonance frequencies or changing the overlap between their electromagnetic fields and make it possible to implement the non-Gaussian operations in cQED. 15 The non-linearities of the cQED devices underlying the non-Gaussian operations are also an exciting avenue to explore. Anharmonic resonators, such as superconducting nonlinear asymmetric inductive element (SNAIL) 32 will be discussed in more detail in later sections of this perspective.
@@ 3.1 Molecular vibronic spectra
Figure 2: Schematic of the superconducting bosonic device setup for simulating the vibration of a water molecule. Here, $\hat{c}_{A}$ and $\hat{c}_{B}$ are the microwave cavity modes representing the two vibrational modes. Transmon qubits $\hat{t}_{A},\hat{t}_{B}$ and readout resonators $\hat{r}_{A},\hat{r}_{B}$ take care of state preparation and measurement. The coupler transmon $\hat{t}_{C}$ is used for Gaussian operations on the cavity modes and the readout resonator $\hat{r}_{c}$ is used for characterization and postselection. Reprinted from Ref. 22.
Molecular spectroscopy is a highly sensitive probe of chemical structure and dynamics. 33, 34 However, the simulation of vibronic spectra, which measures transitions between vibrational states in manifolds that correspond to different electronic states, is a challenging computational task. 35 Specifically, predicting molecular vibronic spectra calls for the efficient calculation of Franck–Condon factors (FCFs), which are given in terms of the overlap of the initial and final vibrational wavefunctions. Recent studies have demonstrated that the calculation of vibronic spectra can be carried out efficiently on bosonic devices. This is because the computation of FCFs is related to the boson sampling problem. 20, 22
To date, cQED devices have been used to demonstrate the vibronic spectra of H2O, O3, NO2, and SO2 molecules.22 In what follows, we discuss the case of water in some detail, as a specific example.
Figure 3: Vibronic spectra for the photoionization of water to the ($\tilde{B}^{2}B_{2}$) excited state of the H2O+ cation, starting from the vibrationless H2O molecule. Reprinted from Ref. 22.
@@
The goal then is to compute the FCFs of the water molecule, which are given by
@@
Here, |$\ket{n,m}$ is the Fock state corresponding to the initial vibronic state, |$\ket{n^{\prime},m^{\prime}}$ is the Fock state corresponding to the final vibronic state, and $\hat{U}_{\text{Dok}}$ is the Doktorov operator representing the Duschinsky transformation in terms of two single-mode displacements, two single-mode squeezing, and one two-mode beam-splitter operations. 22 The Duschinsky transformation encodes the vibronic transition between two vibrational states under a harmonic approximation of the initial and final potential energy surfaces. Thus, after preparing the initial Fock state |$\ket{m,n}$ in the cavity resonators, the Doktorov operator is applied followed by a set of measurements. In the final step, the photon number measurements are averaged to compute the probability of finding the Fock state |$\ket{n^{\prime},m^{\prime}}$, which is equivalent to computing the corresponding FCFs. The measurements can be done by either single-bit extraction, which maps |$\ket{n^{\prime},m^{\prime}}$ to a state of two transmons, or by single-shot photon number resolving detection, which measures the number of photons in a given cavity mode. 22 The simulated vibronic spectra for the photoionization of the water molecule, $\text{H}_{2}\text{O}\xrightarrow{h\nu}\text{H}_{2}\text{O}^{+}+e^{-}$ starting from the vibrational state |$\ket{0,0}$ is shown in Figure 3.
@@ 3.2 Vibrational spectra of nonrigid molecules
The normal mode (harmonic) approximation for molecular vibrations holds when the amplitudes of the oscillations are small. But there is a broad range of nonrigid molecules for which one or more vibrational modes are anharmonic and the associated nuclear displacements are large. Such large amplitude modes are described by complicated potentials that have at least two minima. The best known cases are the bending vibration in the water molecule 36, 37 and the $\nu_{7}$ bending mode of cyanogen iso-thiocyanate (NCNCS) 38, 39.
A successful approach to understanding the complex vibrational spectra of nonrigid molecules consists of approximating their energy potential function with the bottom of a champagne bottle 40. When the excitation energy reaches the local maximum at the origin of the potential, the nature of the vibration switches from bent to linear, and complex features arise in the spectrum. This sudden change in the nature of the excited states of a system as it goes through a critical excitation energy has become known as an excited state quantum phase transition (ESQPT) 41, in analogy with the concept of a quantum phase transition (QPT), which corresponds to a sudden change in the ground state of the system as a control parameter reaches a critical point. 42, 43 One of the main signatures of an ESQPT is a singularity in the density of states at the critical excitation energy, which moves to higher excitation energies as the control parameter increases.
There are numerous systems that exhibit an ESQPT 41 and the first experimental detection of such transition was done in molecules 44, 45, 38, 46, 37, 47, 39. Recently, it has been argued 23 that the superconducting circuit used to generate an anharmonic resonator offers an excellent platform to analyze the spectra of nonrigid molecules and ESQPTs in general. This is because the static effective Hamiltonian of the driven anharmonic resonator 48,
@@
describes a double-well metapotential that captures the ESQPT. In the equation above, $\hat{n}=\hat{a}^{\dagger}\hat{a}$, $K$ is the Kerr nonlinearity and $\xi$ is the control parameter. We denote the eigenvalues of $\hat{H}_{qu}$ by $E$ and the ground state energy by $E_{0}$. Figure 4 shows the main features of the ESQPT associated with $\hat{H}_{qu}$. The left panel displays the excitation energies, $E^{\prime}=(E-E_{0})$ as a function of the control parameter $\xi$. The orange solid line marks the critical energy of the ESQPT for a given value of the control parameter. Below this line, the levels are within the double-well metapotential and pairs of adjacent eigenvalues (black and red lines) merge together (“spectral kissing” 48), while above the ESQPT line the levels are outside the wells and are non-degenerate. The right panel shows the peak of the density of states at the ESQPT energy. This peak diverges in the mean-field limit.
Figure 4: Left: Excitation energies of the Hamiltonian in Eq. (10) as a function of the control parameter $\xi$. Solid black lines are for the levels in the even parity sector, dashed red lines are for the odd parity, and the thick solid orange line marks the energy of the ESQPT. Right: Normalized density of states, where the shade indicates numerical results and the solid line is an analytical curve. The vertical dashed line marks the ESQPT energy (more details in Ref. 23).
The analysis of the spectra of nonrigid molecules with superconducting circuit platforms can be extended to more elaborate scenarios by coupling two or more oscillators. The single-driven nonlinear oscillator examined in Eq. (10) has just one degree of freedom, while a full description of the vibrational bending modes in nonrigid molecules requires two degrees of freedom. By coupling two oscillators, an additional degree of freedom can be accessed, with which molecular spectra including the vibrational angular momentum dependence could be reproduced. A second research line that can be investigated is the use of two coupled qumodes, each one describing a local stretching mode, as a simulator of the transition from normal to local stretching vibrations. In addition to spectroscopy, the superconducting circuit platform discussed here could be used as a simulator for other phenomena related to ESQPTs, such as isomerization reactions 49, 50. The concentration of energy levels at the ESQPT critical point is analogous to the approach of the neighboring energy levels that happen close to an isomerization barrier height. These examples provide new research avenues to explore in the near future.
@@ 3.3 Quantum dynamics
Figure 5: Schematic for seven chromophores in Fenna–Matthews–Olson (FMO) complex, where the arrows represent energy transfer. Reprinted from Ref. 24.
Chemical dynamics is inherently quantum, and for a closed system, it is governed by the time-dependent Schrödinger equation. Within this equation, the generator of dynamics is the molecular Hamiltonian. We have recently shown that the molecular Hamiltonian can be mapped onto a polynomial of bosonic creation and annihilation operators of a single bosonic mode (SBM). 24 Thus, using this so-called SBM mapping allows for the simulation of chemical dynamics on a cQED device.
@@
which is represented in terms of the basis set $\{|\alpha\rangle\}$ of choice, into the following elementary bosonic operators of a single qumode $(\hat{a},\hat{a}^{\dagger})$, as follows:24
@@
where $\hat{\Gamma}_{k}=((k-1)-\hat{n})\hat{a}$. For example, using SBM mapping any $2\times 2$ Hamiltonian such as the Pauli operators can be mapped onto the Hamiltonian of a currently available bosonic quantum device known as driven superconducting nonlinear asymmetric inductive element (SNAIL) 32
$\hat{H}_{\text{SNAIL}}=\hbar\omega\>\hat{a}^{\dagger}\hat{a}+g_{3}\>(\hat{a}+\hat{a}^{\dagger})^{3}~{}~{}.$ (14)
Given that all 1-qubit gates correspond to unitary dynamics generated by a $2\times 2$ Hamiltonian, they can therefore be generated by such SNAIL devices. Furthermore, $4\times 4$ Hamiltonian matrices, which correspond to 2-qubit gates, can also be implemented by combining the SNAIL with a combination of Josephson junctions that gives rise to fourth-order terms in the device Hamiltonian. 24
The SBM mapping has also been used to simulate quantum dynamics in several chemical systems, including energy transfer dynamics in the Fenna–Matthews–Olson (FMO) light-harvesting complex and charge-transfer (redox) chemical reactions within the spin-boson model.24 As an example, we consider the application of SBM mapping to the FMO complex (see Figure 5). Assuming that the energy transfer in FMO occurs across four chlorophyll molecules, 51, 52 a model Hamiltonian for FMO can be represented by the following $4\times 4$ Hermitian matrix (in units of cm-1):51, 52
@@
Figure 6: A schematic protocol for simulating energy transfer in the FMO 4-site model in Eq. (15) with a SNAIL device. Reprinted from Ref. 24.
Figure 7: Comparison between SNAIL simulations with the benchmark data for the population dynamics for the 4-site FMO complex. Reprinted from Ref. 24.
@@
Since many systems of chemical interest are coupled to their surrounding environment, the simulation of open quantum systems on quantum computers has recently gained much attention.53, 14, 54, 55, 56, 57, 58, 59 Most of these simulations are based on the Lindblad equation which relies on the Born–Markov approximation in the system-environment weak coupling limit. 14, 54, 55 Developing quantum algorithms for numerically exact quantum dynamics applied to the regime beyond this weak coupling may involve either doing part of the computations on a classical computer or resorting to more qubits to represent the environmental degrees of freedom. For example, to simulate the dynamics of the spin-boson model, the current qubit-based quantum algorithm developed for the formally exact generalized quantum master equation (GQME) approach 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73 relies on a classical computer to solve the underlying exact memory kernel due to the limited computational resources of current NISQ devices. 14 The hierarchical equations of motion (HEOM) approach is a numerically exact method for open quantum systems,74, 75, 76, 77, 78, 79, 80 which decomposes the environment into a few dissipative bosonic modes, 81, 57 and as a result needs extra qubits for encoding. Bosonic quantum devices provide a natural way to encode the dissipative modes with qumodes to simulate the dynamics of quantum systems coupled to a bosonic environment.

### v3

@@ 1 Introduction
In this paper, we focus on a promising experimental realization of a qumode-based platform that uses circuit quantum electrodynamics (cQED). 15 In this case, superconducting resonators serve as qumodes, with the interactions between resonators controlled via an auxiliary superconducting circuit. 16, 17, 18 Other hardware approaches to bosonic quantum computing include arrays of trapped ions, 19, 20 optical modes, 21 or mechanical resonators. 22
The remainder of this perspective is organized as follows. An introduction to bosonic quantum computing is provided in Section 2. Recent progress in simulating molecular vibronic spectra23, 24 and chemical quantum dynamics 25 on bosonic quantum devices is reviewed in Section 3. The potential for bosonic quantum devices to solve a wide range of timely problems in computational chemistry, including nonadiabatic chemical dynamics, the efficient solution of molecular graph theory problems and electronic structure calculations is highlighted in Section 4. Concluding remarks are provided in Section 5.
@@ 2.1 A primer on bosonic systems
where {|$\{\ket{q}|\>q\in\mathbb{R}\}$ and {|$\{\ket{p}|\>p\in\mathbb{R}\}$ are two possible continuous variable (CV) bases for a single qumode Hilbert space. The transformation between the position and momentum bases is given by the Fourier transform 26
@@
The existence of position and momentum CV bases also makes it possible to represent quantum states in terms of wave functions $\langle q|\psi\rangle=\psi(q)$ and $\langle p|\psi\rangle=\tilde{\psi}(p)$, in the position or momentum representations, respectively. Of particular importance for the bosonic devices under consideration in this paper are Gaussian states, a class of CV states that can be represented in terms of Gaussian functions. 26 One such Gaussian state is the ground state (also known as the vacuum state) of a single qumode, $\langle q|0\rangle=\psi_{0}(q)$. Thus, Gaussian transformations such as phase space displacement and squeezing operators can be defined as operators that map Gaussian states to Gaussian states. 26 Some of the Gaussian unitary operators are listed in Table 1. However, for universal quantum computations, the Gaussian operations have to be combined with non-Gaussian unitaries. 27, 28 The implementation of non-Gaussian operations can be efficiently done on the bosonic circuit quantum electrodynamics (cQED) devices, as discussed below.
@@ 2.2 Hardware
Each resonator can be thought of as a quantum LC circuit, which is a quantized version of a classical LC circuit consisting of an inductor (coil) of inductance $L$ and a capacitor of capacitance $C$. A classical LC circuit behaves as a mechanical system that undergoes simple harmonic motion, such as a mass on a spring. In the circuit, electric charge and current fulfill the role, respectively, of position and velocity in a mechanical spring. The current moves back and forth between the inductor and the capacitor with a resonant frequency given by $\omega=1/\sqrt{LC}$. This is also the frequency of the quantum LC circuit which realizes a quantum harmonic oscillator. The resonator geometry design critically affects the performance of the quantum devices. Early resonator prototypes focused on the planar design because of the convenience of fabricating superconducting metal circuits on the same chip that contains the auxiliary circuits with standard lithographic tools. 29 A range of geometries have been introduced since, including 3D resonators that can produce coherence time up to milliseconds by taking advantage of their hollow structure. 15 The coupling to the auxiliary circuit can now be achieved, for example, by introducing a center pin in a cylindrical resonator, while placing the auxiliary circuit next to the resonator. 18
The auxiliary circuits that are used to couple the qumodes are made of a Josephson junction, a device of two layers of superconducting material separated by a thin insulator. In a superconducting circuit, a Josephson junction substitutes the inductor of the LC circuit and the current is generated by Cooper pairs (pairs of bound electrons) that tunnel through the junction. 30, 31 The Josephson junction becomes insensitive to charge noise in the regime of large Josephson energy to the charging energy ratio, 32 which is commonly referred to as a transmon. 33, 34 The transmon superconducting circuit works as an anharmonic oscillator where the levels are no longer equidistant, which allows for selecting the transitions between energy levels. The two lowest levels are used in the transmon qubits, which are the predominant element in circuit-based quantum hardware, such as the IBM and Google quantum processors. 35, 36
The transmon is coupled to the qumode resonator by extending its superconducting layers to overlap with the resonator’s electromagnetic field. The interaction between the auxiliary circuits and qumodes can be controlled by detuning the qumode and transmon resonance frequencies or changing the overlap between their electromagnetic fields and make it possible to implement the non-Gaussian operations in cQED. 15 The non-linearities of the cQED devices underlying the non-Gaussian operations are also an exciting avenue to explore. Anharmonic resonators, such as superconducting nonlinear asymmetric inductive element (SNAIL) 37 will be discussed in more detail in later sections of this perspective.
@@ 3.1 Molecular vibronic spectra
Figure 2: Schematic of the superconducting bosonic device setup for simulating the vibration of a water molecule. Here, $\hat{c}_{A}$ and $\hat{c}_{B}$ are the microwave cavity modes representing the two vibrational modes. Transmon qubits $\hat{t}_{A},\hat{t}_{B}$ and readout resonators $\hat{r}_{A},\hat{r}_{B}$ take care of state preparation and measurement. The coupler transmon $\hat{t}_{C}$ is used for Gaussian operations on the cavity modes and the readout resonator $\hat{r}_{c}$ is used for characterization and postselection. Reprinted from Ref. 23.
Molecular spectroscopy is a highly sensitive probe of chemical structure and dynamics. 38, 39 However, the simulation of vibronic spectra, which measures transitions between vibrational states in manifolds that correspond to different electronic states, is a challenging computational task. 40 Specifically, predicting molecular vibronic spectra under the Born–Oppenheimer approximation calls for the efficient calculation of Franck–Condon factors (FCFs), which are given in terms of the modulus square of the overlap between the initial and final vibrational wavefunctions corresponding to their potential energy surfaces. Recent studies have demonstrated that the calculation of vibronic spectra can be carried out efficiently on bosonic devices. This is because the computation of FCFs is related to the boson sampling problem. 21, 23, 41
To date, cQED devices have been used to demonstrate the vibronic spectra of H2O, O3, NO2, and SO2 molecules.23 In what follows, we discuss the case of water in some detail, as a specific example.
Figure 3: Vibronic spectra for the photoionization of water to the ($\tilde{B}^{2}B_{2}$) excited state of the H2O+ cation, starting from the vibrationless H2O molecule. Reprinted from Ref. 23.
@@
The computation of FCF involves encoding the relation between initial and final normal coordinates corresponding to the potential energy surfaces for the vibronic transitions, which is known as the Duschinsky transformation. 42 In the Hilbert space representation, the Duschinsky transformation can be represented by the so-called Doktorov operator, 43, 44 and the FCF for two vibrational modes becomes
@@
Here, |$\ket{n,m}$ is the Fock state corresponding to the initial vibronic state, |$\ket{n^{\prime},m^{\prime}}$ is the Fock state corresponding to the final vibronic state, and $\hat{U}_{\text{Dok}}$ is the Doktorov operator representing the Duschinsky transformation in terms of two single-mode displacement, two single-mode squeezing, and one two-mode beam-splitter operations, which correspond to the displacement, distortion, and rotation of the potential energy surfaces due to the vibronic transition. 21, 23, 41 Thus, after preparing the initial Fock state |$\ket{m,n}$ in the cavity resonators, the Doktorov operator is applied followed by a set of measurements. In the final step, the photon number measurements are averaged to compute the probability of finding the Fock state |$\ket{n^{\prime},m^{\prime}}$, which is equivalent to computing the corresponding FCFs. The measurements can be done by either single-bit extraction, which maps |$\ket{n^{\prime},m^{\prime}}$ to a state of two transmons, or by single-shot photon number resolving detection, which measures the number of photons in a given cavity mode. 23 The simulated vibronic spectra for the photoionization of the water molecule, $\text{H}_{2}\text{O}\xrightarrow{h\nu}\text{H}_{2}\text{O}^{+}+e^{-}$ starting from the vibrational state |$\ket{0,0}$ is shown in Figure 3. It has been shown that simulating the same spectra needs 8 qubits and $\mathcal{O}(10^{3})$ gates for a qubit-based algorithm, 23 validating the advantage of bosonic quantum devices for bosonic problems.
@@ 3.2 Vibrational spectra of nonrigid molecules
The normal mode (harmonic) approximation for molecular vibrations holds when the amplitudes of the oscillations are small. But there is a broad range of nonrigid molecules for which one or more vibrational modes are anharmonic and the associated nuclear displacements are large. For example, such large amplitude modes are described by complicated potentials that have at least two minima. The best known cases are the bending vibration in the water molecule 45, 46 and the $\nu_{7}$ bending mode of cyanogen iso-thiocyanate (NCNCS) 47, 48.
A successful approach to understanding the complex vibrational spectra of nonrigid molecules consists of approximating their energy potential function as discussed in Ref. 49. When the excitation energy reaches the local maximum at the origin of the potential, the nature of the vibration switches from bent to linear, and complex features arise in the spectrum. This sudden change in the nature of the excited states of a system as it goes through a critical excitation energy has become known as an excited state quantum phase transition (ESQPT) 50, in analogy with the concept of a quantum phase transition (QPT), which corresponds to a sudden change in the ground state of the system as a control parameter reaches a critical point. 51, 52 One of the main signatures of an ESQPT is a singularity in the density of states at the critical excitation energy, which moves to higher excitation energies as the control parameter increases.
There are numerous systems that exhibit an ESQPT 50 and the first experimental detection of such transition was done in molecules 53, 54, 47, 55, 46, 56, 48. Recently, it has been argued 24 that the superconducting circuit used to generate an anharmonic resonator offers an excellent platform to analyze the spectra of nonrigid molecules and ESQPTs in general. This is because the static effective Hamiltonian of the driven anharmonic resonator 57,
@@
describes a double-well metapotential that captures the ESQPT. In the equation above, $\hat{n}=\hat{a}^{\dagger}\hat{a}$, $K$ is the Kerr nonlinearity and $\xi$ is the control parameter. We denote the eigenvalues of $\hat{H}_{qu}$ by $E$ and the ground state energy by $E_{0}$. Figure 4 shows the main features of the ESQPT associated with $\hat{H}_{qu}$. The left panel displays the excitation energies, $E^{\prime}=(E-E_{0})$ as a function of the control parameter $\xi$. The orange solid line marks the critical energy of the ESQPT for a given value of the control parameter. Below this line, the levels are within the double-well metapotential and pairs of adjacent eigenvalues (black and red lines) merge together (“spectral kissing” 57), while above the ESQPT line the levels are outside the wells and are non-degenerate. The right panel shows the peak of the density of states at the ESQPT energy. This peak diverges in the mean-field limit.
Figure 4: Left: Excitation energies of the Hamiltonian in Eq. (10) as a function of the control parameter $\xi$. Solid black lines are for the levels in the even parity sector, dashed red lines are for the odd parity, and the thick solid orange line marks the energy of the ESQPT. Right: Normalized density of states, where the shade indicates numerical results and the solid line is an analytical curve. The vertical dashed line marks the ESQPT energy (more details in Ref. 24).
The analysis of the spectra of nonrigid molecules with superconducting circuit platforms can be extended to more elaborate scenarios by coupling two or more oscillators. The single-driven nonlinear oscillator examined in Eq. (10) has just one degree of freedom, while a full description of the vibrational bending modes in nonrigid molecules requires two degrees of freedom. By coupling two oscillators, an additional degree of freedom can be accessed, with which molecular spectra including the vibrational angular momentum dependence could be reproduced. A second research line that can be investigated is the use of two coupled qumodes, each one describing a local stretching mode, as a simulator of the transition from normal to local stretching vibrations. In addition to spectroscopy, the superconducting circuit platform discussed here could be used as a simulator for other phenomena related to ESQPTs, such as isomerization reactions 58, 59. The concentration of energy levels at the ESQPT critical point is analogous to the approach of the neighboring energy levels that happen close to an isomerization barrier height. These examples provide new research avenues to explore in the near future.
@@ 3.3 Quantum dynamics
Figure 5: Schematic for seven chromophores in Fenna–Matthews–Olson (FMO) complex, where the arrows represent energy transfer. Reprinted from Ref. 25.
Chemical dynamics is inherently quantum, and for a closed system, it is governed by the time-dependent Schrödinger equation. Within this equation, the generator of dynamics is the molecular Hamiltonian. It has recently been shown that the molecular Hamiltonian can be mapped onto a polynomial of bosonic creation and annihilation operators of a single bosonic mode (SBM). 25 Thus, using this so-called SBM mapping allows for the simulation of chemical dynamics on a cQED device.
@@
which is represented in terms of the basis set $\{|\alpha\rangle\}$ of choice, into the following elementary bosonic operators of a single qumode $(\hat{a},\hat{a}^{\dagger})$, as follows:25
@@
where $\hat{\Gamma}_{k}=((k-1)-\hat{n})\hat{a}$. For example, the Pauli-Z matrix can be represented as
$Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}\mapsto 1-2\>\hat{a}^{\dagger}\hat{a},$ (14)
using the SBM mapping. Similarly, any $2\times 2$ Hamiltonian such as the Pauli operators can be mapped onto the Hamiltonian of a currently available bosonic quantum device known as driven superconducting nonlinear asymmetric inductive element (SNAIL) 37, 60, 61
$\hat{H}_{\text{SNAIL}}=\hbar\omega\>\hat{a}^{\dagger}\hat{a}+g_{3}\>(\hat{a}+\hat{a}^{\dagger})^{3}.$ (15)
Given that all 1-qubit gates correspond to unitary dynamics generated by a $2\times 2$ Hamiltonian, they can therefore be generated by such SNAIL devices. Furthermore, $4\times 4$ Hamiltonian matrices, which correspond to 2-qubit gates, can also be implemented by combining the SNAIL with a combination of Josephson junctions that gives rise to fourth-order terms in the device Hamiltonian. 25
The SBM mapping has also been used to simulate quantum dynamics in several chemical systems, including energy transfer dynamics in the Fenna–Matthews–Olson (FMO) light-harvesting complex and charge-transfer (redox) chemical reactions within the spin-boson model.25 As an example, we consider the application of SBM mapping to the FMO complex (see Figure 5). Assuming that the energy transfer in FMO occurs across four chlorophyll molecules, 62, 63 a model Hamiltonian for FMO can be represented by the following $4\times 4$ Hermitian matrix (in units of cm-1):62, 63
@@
Figure 6: A schematic protocol for simulating energy transfer in the FMO 4-site model in Eq. (16) with a SNAIL device. Adapted from Ref. 25.
Figure 7: Comparison between SNAIL simulations with the benchmark data for the population dynamics for the 4-site FMO complex. Reprinted from Ref. 25.
@@
Since many systems of chemical interest are coupled to their surrounding environment, the simulation of open quantum systems on quantum computers has recently gained much attention.64, 14, 65, 66, 67, 68, 69, 70 Most of these simulations are based on the Lindblad equation which relies on the Born–Markov approximation in the system-environment weak coupling limit. 14, 65, 66 Developing quantum algorithms for numerically exact quantum dynamics applied to the regime beyond this weak coupling may involve either doing part of the computations on a classical computer or resorting to more qubits to represent the environmental degrees of freedom. For example, to simulate the dynamics of the spin-boson model, the current qubit-based quantum algorithm developed for the formally exact generalized quantum master equation (GQME) approach 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84 relies on a classical computer to solve the underlying exact memory kernel due to the limited computational resources of current NISQ devices. 14 The hierarchical equations of motion (HEOM) approach is a numerically exact method for open quantum systems,85, 86, 87, 88, 89, 90, 91 which decomposes the environment into a few dissipative bosonic modes, 92, 68 and as a result needs extra qubits for encoding. Bosonic quantum devices provide a natural way to encode the dissipative modes with qumodes to simulate the dynamics of quantum systems coupled to a bosonic environment.